# Climb: Class-imbalanced Learning Benchmark on Tabular Data

### v1

@@ A.3 Additional Results
RUS CC IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SPE BC BRF EE UBS UBA OBS SMBS OBA SMBA CS AdaC AdaBS AsyBS
bwin_amlb 6.8 3.8 15.2 588.0 4.0 5.0 7.4 30.1 13.5 14.2 8.6 4.2 5.4 7.4 12.4 8.4 331.1 339.6 182.7 309.5 391.7 324.9 562.5 650.8 424.3 574.9 4.0 447.4 394.1 439.9
mozilla4 34.0 21.6 1445.1 4683.8 59.2 72.0 137.2 490.0 111.9 180.9 159.1 35.6 57.4 81.4 1724.8 89.2 2066.3 1973.1 1135.8 1926.5 2744.7 1912.8 4319.8 4665.6 3185.8 4464.1 35.8 4039.8 3804.6 4033.1
mc2 1.8 1.4 22.8 412.0 2.0 2.8 4.0 6.4 4.0 4.0 4.5 2.4 4.0 3.8 7.0 3.6 289.4 251.3 153.0 223.2 233.9 212.7 468.7 384.5 249.8 314.0 2.2 216.0 237.8 218.1
wholesale-customers 2.2 2.0 15.6 483.0 3.0 4.3 5.3 13.8 21.1 8.8 8.5 2.2 3.6 5.6 7.7 4.6 234.7 227.4 152.7 210.1 205.7 217.8 314.0 340.4 233.9 339.2 1.8 195.6 192.3 195.2
vertebra-column 1.0 1.6 9.0 426.3 2.4 2.4 4.0 12.8 5.4 6.2 5.8 1.4 3.0 3.8 5.8 3.4 202.1 186.8 140.4 203.9 219.8 202.7 255.0 302.3 218.4 299.7 1.4 154.7 170.5 161.0
law-school-admission-bianry 39.7 29.2 2989.7 4231.6 213.1 362.9 324.3 4902.8 1903.1 603.7 852.0 51.3 198.6 429.6 5917.3 418.7 3513.4 3640.4 1217.3 3264.7 3113.1 3240.3 6999.0 12094.0 5698.5 11381.4 49.2 5006.4 5186.8 5129.1
bank32nh 201.5 107.4 807.1 9568.5 159.8 326.8 245.5 479.5 950.3 361.0 335.8 235.0 318.1 343.0 969.8 362.1 9159.3 9385.4 5544.9 9501.8 20479.8 9490.1 26646.9 33913.7 18378.3 23142.3 211.1 22409.9 21346.6 22636.8
elevators 74.8 47.2 2234.0 5836.4 155.6 541.3 501.2 3050.8 838.7 624.0 689.1 85.5 180.4 312.7 1821.2 333.0 4227.1 4213.5 2390.0 4211.4 6506.5 4222.4 9354.2 14430.7 7879.3 14603.9 68.3 7632.1 7224.1 7592.4
cpu_small 43.4 25.4 583.3 3747.3 111.6 204.8 233.6 1068.4 645.8 325.7 286.1 53.3 107.8 167.9 370.9 186.4 2490.6 2405.5 1278.5 2335.3 3501.5 2358.9 6015.1 8119.0 4390.6 6527.4 47.6 5012.5 4770.8 4962.3
Credit_Approval_Classification 5.4 3.6 37.2 630.0 5.0 7.9 6.8 7.6 9.8 11.6 22.6 8.0 11.0 12.8 26.2 10.5 418.6 416.7 196.7 394.4 489.3 378.4 749.1 1012.2 616.6 810.7 4.0 489.0 542.5 532.1
house_8L 132.3 76.7 3239.4 9850.5 372.2 767.4 1028.1 3412.6 5327.3 3218.4 2836.1 157.0 484.7 1369.7 9715.4 1406.9 12746.5 14529.8 5977.8 6511.0 15241.0 8826.1 48039.0 54452.8 20050.6 18944.1 137.2 15453.1 15462.0 15755.9
house_16H 292.0 171.9 4210.7 19484.0 362.0 1181.5 972.5 3529.1 1794.5 1363.3 1282.7 350.1 541.3 732.6 3846.4 816.0 13168.7 15406.8 9104.9 14308.7 23041.9 14226.6 39188.7 51926.8 27528.0 40540.8 311.8 32000.0 32086.1 32211.3
phoneme 18.0 12.0 238.6 2420.8 21.6 27.7 26.7 76.2 43.9 68.2 62.6 22.3 34.6 38.4 243.7 46.7 1067.2 1088.1 659.2 1060.6 1449.9 1046.8 2619.8 3332.5 1772.9 2289.4 21.0 2159.5 2079.5 2141.4
ilpd-numeric 3.2 2.1 16.9 571.4 3.4 4.4 5.7 5.6 12.2 10.2 10.2 3.0 5.2 7.0 12.8 7.8 286.3 257.1 166.0 264.3 281.9 258.1 430.9 536.5 345.7 471.0 2.4 286.0 322.1 303.6
planning-relax 1.6 1.6 5.4 433.5 3.0 3.2 3.6 3.4 4.4 5.4 5.6 2.0 5.0 5.1 6.0 4.0 233.6 211.7 137.8 196.2 207.6 195.3 297.5 336.5 224.8 308.1 1.8 180.1 195.6 183.2
MiniBooNE 11421.9 5354.3 352768.8 350197.9 13489.4 50621.8 38412.4 213900.5 83752.8 58883.8 50310.3 14874.9 21436.0 30543.4 191814.8 31782.7 379865.7 452074.4 44202.2 481420.7 846681.6 486466.0 1598043.1 2183156.2 1132603.5 1665075.7 11828.5 1245098.1 1213771.8 1221899.0
machine_cpu 0.8 1.2 5.6 379.9 2.0 2.2 2.4 4.2 4.4 4.6 5.0 1.0 1.8 3.6 4.4 3.0 186.3 164.7 124.1 168.0 165.6 167.8 196.7 184.3 170.6 250.7 1.0 108.4 104.8 101.7
telco-customer-churn 24.8 13.8 503.9 1702.6 43.1 124.6 97.6 1044.6 738.6 158.1 170.9 33.1 56.4 91.3 571.5 81.1 1972.4 1909.0 557.8 1660.9 1386.2 1633.4 3187.8 4633.4 3582.0 4598.3 26.0 2168.9 2199.8 2185.7
haberman 1.0 1.4 6.8 448.2 2.4 2.4 4.2 16.1 22.6 6.4 5.4 1.4 2.2 4.0 6.4 3.8 224.5 183.6 133.3 172.2 176.8 176.0 234.5 227.2 185.7 270.4 1.0 116.8 118.9 127.8
vehicle 3.0 2.8 20.4 582.0 3.9 5.4 9.0 22.9 9.0 7.9 11.0 4.4 7.2 8.2 12.3 10.3 324.0 295.0 178.8 277.7 323.3 280.4 522.5 842.9 405.6 628.2 2.8 323.9 329.8 324.7
cpu 1.0 1.0 5.4 382.7 2.2 1.8 3.2 3.8 3.8 3.1 4.3 1.2 1.8 2.4 4.0 2.8 217.4 191.7 130.9 180.8 168.6 178.6 204.3 196.1 224.2 259.3 1.0 100.1 107.1 102.5
ada 11.4 7.7 229.7 1107.1 17.4 54.2 60.7 287.1 143.9 64.2 79.4 23.3 30.4 38.7 185.7 44.2 815.7 787.2 360.2 823.9 934.2 817.1 2086.4 2975.8 1769.1 2533.7 13.0 1320.6 1366.2 1374.3
adult 451.7 200.2 72817.1 14514.1 1781.1 8740.8 6973.0 35287.1 16080.4 10662.0 9447.5 800.5 1419.8 3167.7 50721.2 3400.3 28388.0 27708.0 6958.4 25068.7 44634.6 30450.9 242023.7 297819.9 128392.3 217251.2 445.7 149107.8 142462.4 98830.2
blood-transfusion-service-center 1.2 1.2 17.0 519.7 2.5 3.4 2.6 32.7 6.2 9.4 10.6 2.2 3.6 4.8 15.2 4.6 239.9 232.7 154.3 209.7 221.8 226.9 326.5 386.1 237.0 332.9 1.2 181.1 184.4 189.3
default-of-credit-card-clients 437.4 150.2 11897.6 20779.0 499.9 2005.0 1574.4 5697.7 3077.9 2295.9 2156.9 492.7 736.7 1063.4 15399.6 1027.3 12880.0 26953.1 7187.4 26209.6 24258.7 14772.1 58346.9 82664.8 70583.7 81441.7 394.7 50115.4 46239.5 49296.7
Customer_Churn_Classification 1835.5 784.6 466783.2 93533.7 10638.3 60348.6 48486.9 337850.7 115694.5 70833.6 64674.9 2640.4 7218.6 19188.1 563263.8 19161.5 67007.0 70532.8 39838.3 68900.0 91764.8 68767.4 282308.1 723570.9 225611.1 565491.7 1674.5 233768.6 245187.2 234979.4
SPECTF 2.8 2.0 6.6 454.3 2.6 3.6 3.7 7.6 4.4 4.8 6.4 4.2 9.1 9.8 11.6 9.4 255.5 229.0 165.8 226.6 280.6 238.4 526.9 916.9 368.9 531.3 2.8 303.5 319.5 307.1
Medical-Appointment-No-Shows 500.7 206.2 159874.0 33280.6 4431.4 24476.0 19701.5 71004.8 42701.5 28274.0 24966.6 830.3 2603.7 7738.6 855622.3 7973.6 27234.0 29654.6 10958.9 25832.4 20695.4 27527.3 82677.8 137588.1 84576.7 195401.4 549.1 50300.9 49717.2 51772.3
JapaneseVowels 68.0 23.2 511.6 4572.5 130.6 220.0 464.5 1468.1 654.1 375.3 422.4 105.1 217.7 313.5 370.0 288.0 2156.2 2337.2 988.1 2354.9 2845.0 2385.2 11809.2 23292.2 7908.2 12345.3 83.1 8072.1 8403.0 8069.3
ibm-employee-attrition 11.4 4.4 35.2 841.8 6.6 17.8 27.2 44.6 30.5 20.2 28.4 16.6 27.3 30.9 53.0 26.2 634.9 615.6 311.2 541.1 602.7 529.5 1544.5 2578.5 1444.5 1962.3 11.4 1199.1 1067.7 1165.2
first-order-theorem-proving 260.6 123.0 1076.6 14412.8 189.2 564.0 376.0 377.1 451.6 581.2 696.8 494.5 1209.6 1387.6 16399.4 1393.4 14029.4 14742.7 2543.1 14821.9 22683.2 15135.4 54801.6 149905.1 60861.9 113144.1 137.9 63868.9 70784.0 67082.8
user-knowledge 1.2 1.0 9.4 445.3 3.2 2.4 7.0 6.0 3.6 9.6 8.0 1.6 5.2 8.0 20.8 1.2 249.5 182.0 136.8 174.6 186.5 168.1 250.2 284.3 205.6 496.7 1.4 139.2 153.2 137.8
online-shoppers-intention 59.0 35.7 932.5 3352.5 70.6 357.1 345.9 1500.2 1245.3 392.7 432.3 89.4 152.9 191.9 1065.0 193.9 2719.6 3234.2 1051.6 2275.9 2352.1 2280.6 8023.8 15974.0 7939.9 11193.9 57.8 5435.3 5324.7 5426.7
kc1 11.0 4.8 43.4 1018.0 4.2 17.2 28.8 145.8 145.9 21.4 29.4 12.6 27.2 28.6 62.2 29.0 548.6 559.3 283.7 444.9 551.6 467.4 2477.4 4628.3 1163.0 1747.6 9.4 1503.6 1808.2 1488.7
thoracic-surgery 1.4 1.0 7.0 462.5 2.0 2.6 2.8 9.0 17.9 3.8 7.6 2.2 3.2 4.0 9.3 4.0 222.5 200.9 134.8 199.3 253.7 211.5 348.6 415.7 265.5 358.0 1.8 192.2 200.4 200.9
UCI_churn 22.7 6.2 77.9 1875.7 11.6 46.2 39.6 99.2 70.3 47.2 64.7 25.6 41.8 45.8 106.6 46.0 688.5 657.5 342.2 705.0 981.7 687.2 2986.1 4644.7 2223.0 3048.3 24.7 2640.1 2433.5 2638.2
arsenic-female-bladder 1.0 1.0 7.0 529.0 2.0 2.4 2.4 8.6 9.1 9.4 9.5 2.4 4.8 5.0 11.3 4.6 249.8 199.8 154.9 196.6 209.8 186.1 287.5 321.7 217.3 314.1 1.4 153.1 152.8 157.3
okcupid_stem 677.7 170.8 33141.2 29727.6 1770.0 13170.7 9867.2 9610.6 9452.7 12852.7 20467.5 3914.3 5342.9 8847.6 319856.7 11635.3 60835.2 60473.4 5368.4 43825.6 39398.3 46729.0 223807.9 215394.0 349760.3 480048.2 278.4 98005.7 102574.4 99067.5
ecoli 0.8 1.2 7.8 396.5 3.6 2.2 6.6 17.5 9.4 0.8 7.2 1.6 4.8 7.6 17.6 8.6 243.7 179.0 137.0 172.7 177.2 179.7 252.0 292.5 206.6 506.3 1.4 133.2 139.4 129.0
pc4 8.6 4.0 29.2 887.2 6.8 15.4 22.0 100.6 298.3 64.6 170.8 20.4 153.4 106.4 198.1 201.8 936.3 1917.4 1346.5 1291.6 1792.9 906.4 5024.9 14518.1 2946.5 2456.0 10.1 984.5 1078.9 975.6
bank-marketing 22.6 5.9 140.6 1419.9 17.8 71.1 70.0 285.6 250.8 80.3 104.3 36.4 70.0 84.8 178.2 70.3 1097.8 1163.9 311.1 875.8 746.8 857.2 3243.3 6964.4 3621.8 4904.8 22.4 2288.0 2016.9 2198.1
Diabetes-130-Hospitals_(Fairlearn) 561.7 142.8 106332.6 26681.7 2729.1 23801.1 20950.1 120267.6 50427.3 26071.4 24390.4 1099.3 2066.7 4550.6 3751167.6 4560.9 22111.3 22994.3 6756.0 20215.9 14156.5 20308.6 102571.3 177673.0 115777.3 160993.5 577.7 51156.7 53470.3 50926.5
Otto-Group-Product-Classification-Challenge 911.5 210.1 12390.7 27696.4 1530.4 12848.0 38053.7 53203.7 51895.8 50469.0 57542.1 2118.4 21850.3 58771.5 1403163.3 58822.2 66475.9 73756.5 12228.7 23946.1 26298.5 23410.6 246684.7 1259334.9 199213.6 1077151.2 911.8 94282.3 89978.2 94552.6
eucalyptus 16.0 4.7 88.7 1424.1 9.8 52.1 45.9 137.7 72.9 63.0 92.6 26.4 47.2 71.4 692.7 70.0 866.6 735.0 282.9 590.3 563.3 560.5 2452.2 3839.1 2338.2 3219.2 17.2 1578.1 1598.7 1574.5
pendigits 24.0 8.0 463.6 1996.3 36.1 228.1 217.9 816.1 579.0 194.6 311.8 51.2 146.3 161.4 205.6 173.0 883.3 821.5 505.2 865.2 905.5 879.9 5336.6 17508.9 4429.2 9634.1 22.2 2569.5 2244.1 2550.7
pc3 18.2 3.4 26.0 1001.1 5.4 20.5 16.8 35.7 28.5 23.1 34.3 24.2 52.8 57.2 59.8 49.2 500.7 456.7 258.8 444.8 475.9 444.1 2385.4 5365.0 1605.7 2869.2 13.8 1425.5 1567.1 1397.1
page-blocks-bin 24.2 4.8 125.4 2337.0 19.8 60.8 114.2 263.1 166.3 104.8 112.7 21.4 42.5 55.6 107.4 53.4 521.9 517.8 274.6 477.5 617.5 492.6 2466.9 4182.5 1742.8 2525.9 19.0 2289.6 1803.9 2281.3
optdigits 30.3 7.8 239.6 1410.9 18.8 111.6 136.8 333.9 372.0 111.4 159.5 60.6 184.4 266.6 239.9 247.2 879.5 874.6 390.8 995.2 1176.1 969.3 6458.1 21344.9 5498.7 12072.3 28.0 3027.8 2535.2 2986.0
mfeat-karhunen 64.0 7.8 51.2 2307.1 11.0 74.7 81.8 121.1 104.4 86.5 105.6 61.2 118.0 124.3 160.7 161.4 913.3 847.1 368.1 809.2 2082.6 827.9 6111.3 13804.2 4229.5 6882.7 47.6 7206.2 5055.5 7092.1
[…]

### v2

@@ C.4 Comparison with BAF Benchmark
RUS $C$ IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SENN STom SPE BC BRF $E$ UBS UBA OBS SMBS OBA SMBA AdaC AdaBS AsyBS CS XGB LGB CAT
bwin_amlb 6.8 3.8 15.2 588.0 4.0 5.0 7.4 30.1 13.5 14.2 8.6 4.2 5.4 7.4 12.4 8.4 33.4 23.2 331.1 339.6 182.7 309.5 391.7 324.9 562.5 650.8 424.3 574.9 447.4 394.1 439.9 4.0 272.5 74.0 133.1
mozilla4 34.0 21.6 1445.1 4683.8 59.2 72.0 137.2 490.0 111.9 180.9 159.1 35.6 57.4 81.4 1724.8 89.2 204.9 211.9 2066.3 1973.1 1135.8 1926.5 2744.7 1912.8 4319.8 4665.6 3185.8 4464.1 4039.8 3804.6 4033.1 35.8 264.1 184.7 220.0
mc2 1.8 1.4 22.8 412.0 2.0 2.8 4.0 6.4 4.0 4.0 4.5 2.4 4.0 3.8 7.0 3.6 17.1 7.5 289.4 251.3 153.0 223.2 233.9 212.7 468.7 384.5 249.8 314.0 216.0 237.8 218.1 2.2 500.3 167.3 234.9
wholesale-customers 2.2 2.0 15.6 483.0 3.0 4.3 5.3 13.8 21.1 8.8 8.5 2.2 3.6 5.6 7.7 4.6 32.1 19.7 234.7 227.4 152.7 210.1 205.7 217.8 314.0 340.4 233.9 339.2 195.6 192.3 195.2 1.8 413.7 1697.6 157.1
vertebra-column 1.0 1.6 9.0 426.3 2.4 2.4 4.0 12.8 5.4 6.2 5.8 1.4 3.0 3.8 5.8 3.4 31.5 18.6 202.1 186.8 140.4 203.9 219.8 202.7 255.0 302.3 218.4 299.7 154.7 170.5 161.0 1.4 237.2 56.8 105.1
law-school-admission-bianry 39.7 29.2 2989.7 4231.6 213.1 362.9 324.3 4902.8 1903.1 603.7 852.0 51.3 198.6 429.6 5917.3 418.7 1376.1 975.9 3513.4 3640.4 1217.3 3264.7 3113.1 3240.3 6999.0 12094.0 5698.5 11381.4 5006.4 5186.8 5129.1 49.2 267.1 194.0 233.3
bank32nh 201.5 107.4 807.1 9568.5 159.8 326.8 245.5 479.5 950.3 361.0 335.8 235.0 318.1 343.0 969.8 362.1 340.3 531.1 9159.3 9385.4 5544.9 9501.8 20479.8 9490.1 26646.9 33913.7 18378.3 23142.3 22409.9 21346.6 22636.8 211.1 352.4 187.3 324.3
elevators 74.8 47.2 2234.0 5836.4 155.6 541.3 501.2 3050.8 838.7 624.0 689.1 85.5 180.4 312.7 1821.2 333.0 284.2 326.5 4227.1 4213.5 2390.0 4211.4 6506.5 4222.4 9354.2 14430.7 7879.3 14603.9 7632.1 7224.1 7592.4 68.3 288.5 247.0 256.6
cpu_small 43.4 25.4 583.3 3747.3 111.6 204.8 233.6 1068.4 645.8 325.7 286.1 53.3 107.8 167.9 370.9 186.4 754.4 547.3 2490.6 2405.5 1278.5 2335.3 3501.5 2358.9 6015.1 8119.0 4390.6 6527.4 5012.5 4770.8 4962.3 47.6 273.6 173.0 184.8
Credit_Approval_Classification 5.4 3.6 37.2 630.0 5.0 7.9 6.8 7.6 9.8 11.6 22.6 8.0 11.0 12.8 26.2 10.5 33.2 29.8 418.6 416.7 196.7 394.4 489.3 378.4 749.1 1012.2 616.6 810.7 489.0 542.5 532.1 4.0 325.9 288.6 220.9
house_8L 132.3 76.7 3239.4 9850.5 372.2 767.4 1028.1 3412.6 5327.3 3218.4 2836.1 157.0 484.7 1369.7 9715.4 1406.9 3566.2 2373.0 12746.5 14529.8 5977.8 6511.0 15241.0 8826.1 48039.0 54452.8 20050.6 18944.1 15453.1 15462.0 15755.9 137.2 287.3 205.9 251.7
house_16H 292.0 171.9 4210.7 19484.0 362.0 1181.5 972.5 3529.1 1794.5 1363.3 1282.7 350.1 541.3 732.6 3846.4 816.0 735.2 927.1 13168.7 15406.8 9104.9 14308.7 23041.9 14226.6 39188.7 51926.8 27528.0 40540.8 32000.0 32086.1 32211.3 311.8 322.6 186.6 290.7
phoneme 18.0 12.0 238.6 2420.8 21.6 27.7 26.7 76.2 43.9 68.2 62.6 22.3 34.6 38.4 243.7 46.7 107.4 96.6 1067.2 1088.1 659.2 1060.6 1449.9 1046.8 2619.8 3332.5 1772.9 2289.4 2159.5 2079.5 2141.4 21.0 254.0 165.1 145.0
ilpd-numeric 3.2 2.1 16.9 571.4 3.4 4.4 5.7 5.6 12.2 10.2 10.2 3.0 5.2 7.0 12.8 7.8 33.5 22.2 286.3 257.1 166.0 264.3 281.9 258.1 430.9 536.5 345.7 471.0 286.0 322.1 303.6 2.4 268.3 83.9 138.4
planning-relax 1.6 1.6 5.4 433.5 3.0 3.2 3.6 3.4 4.4 5.4 5.6 2.0 5.0 5.1 6.0 4.0 33.4 19.1 233.6 211.7 137.8 196.2 207.6 195.3 297.5 336.5 224.8 308.1 180.1 195.6 183.2 1.8 342.5 22.7 81.7
MiniBooNE 11421.9 5354.3 352768.8 350197.9 13489.4 50621.8 38412.4 213900.5 83752.8 58883.8 50310.3 14874.9 21436.0 30543.4 191814.8 31782.7 22782.6 32378.0 379865.7 452074.4 44202.2 481420.7 846681.6 486466.0 1598043.1 2183156.2 1132603.5 1665075.7 1245098.1 1213771.8 1221899.0 11828.5 700.9 438.5 1566.4
machine_cpu 0.8 1.2 5.6 379.9 2.0 2.2 2.4 4.2 4.4 4.6 5.0 1.0 1.8 3.6 4.4 3.0 30.8 17.3 186.3 164.7 124.1 168.0 165.6 167.8 196.7 184.3 170.6 250.7 108.4 104.8 101.7 1.0 284.1 35.6 70.1
telco-customer-churn 24.8 13.8 503.9 1702.6 43.1 124.6 97.6 1044.6 738.6 158.1 170.9 33.1 56.4 91.3 571.5 81.1 115.6 134.5 1972.4 1909.0 557.8 1660.9 1386.2 1633.4 3187.8 4633.4 3582.0 4598.3 2168.9 2199.8 2185.7 26.0 265.3 213.2 315.3
haberman 1.0 1.4 6.8 448.2 2.4 2.4 4.2 16.1 22.6 6.4 5.4 1.4 2.2 4.0 6.4 3.8 30.9 17.8 224.5 183.6 133.3 172.2 176.8 176.0 234.5 227.2 185.7 270.4 116.8 118.9 127.8 1.0 376.8 236.1 88.1
vehicle 3.0 2.8 20.4 582.0 3.9 5.4 9.0 22.9 9.0 7.9 11.0 4.4 7.2 8.2 12.3 10.3 26.7 25.5 324.0 295.0 178.8 277.7 323.3 280.4 522.5 842.9 405.6 628.2 323.9 329.8 324.7 2.8 251.8 111.4 145.4
cpu 1.0 1.0 5.4 382.7 2.2 1.8 3.2 3.8 3.8 3.1 4.3 1.2 1.8 2.4 4.0 2.8 6.2 5.5 217.4 191.7 130.9 180.8 168.6 178.6 204.3 196.1 224.2 259.3 100.1 107.1 102.5 1.0 339.0 137.8 138.2
ada 11.4 7.7 229.7 1107.1 17.4 54.2 60.7 287.1 143.9 64.2 79.4 23.3 30.4 38.7 185.7 44.2 92.5 104.1 815.7 787.2 360.2 823.9 934.2 817.1 2086.4 2975.8 1769.1 2533.7 1320.6 1366.2 1374.3 13.0 254.9 222.7 224.6
adult 451.7 200.2 72817.1 14514.1 1781.1 8740.8 6973.0 35287.1 16080.4 10662.0 9447.5 800.5 1419.8 3167.7 50721.2 3400.3 2229.0 2207.9 28388.0 27708.0 6958.4 25068.7 44634.6 30450.9 242023.7 297819.9 128392.3 217251.2 149107.8 142462.4 98830.2 445.7 481.8 379.7 1031.1
blood-transfusion-service-center 1.2 1.2 17.0 519.7 2.5 3.4 2.6 32.7 6.2 9.4 10.6 2.2 3.6 4.8 15.2 4.6 31.8 19.4 239.9 232.7 154.3 209.7 221.8 226.9 326.5 386.1 237.0 332.9 181.1 184.4 189.3 1.2 259.1 86.5 89.2
default-of-credit-card-clients 437.4 150.2 11897.6 20779.0 499.9 2005.0 1574.4 5697.7 3077.9 2295.9 2156.9 492.7 736.7 1063.4 15399.6 1027.3 1043.4 1350.3 12880.0 26953.1 7187.4 26209.6 24258.7 14772.1 58346.9 82664.8 70583.7 81441.7 50115.4 46239.5 49296.7 394.7 320.1 196.1 530.7
Customer_Churn_Classification 1835.5 784.6 466783.2 93533.7 10638.3 60348.6 48486.9 337850.7 115694.5 70833.6 64674.9 2640.4 7218.6 19188.1 563263.8 19161.5 12326.0 13801.1 67007.0 70532.8 39838.3 68900.0 91764.8 68767.4 282308.1 723570.9 225611.1 565491.7 233768.6 245187.2 234979.4 1674.5 485.7 488.9 1241.7
SPECTF 2.8 2.0 6.6 454.3 2.6 3.6 3.7 7.6 4.4 4.8 6.4 4.2 9.1 9.8 11.6 9.4 17.6 21.4 255.5 229.0 165.8 226.6 280.6 238.4 526.9 916.9 368.9 531.3 303.5 319.5 307.1 2.8 241.0 201.9 207.2
Medical-Appointment-No-Shows 500.7 206.2 159874.0 33280.6 4431.4 24476.0 19701.5 71004.8 42701.5 28274.0 24966.6 830.3 2603.7 7738.6 855622.3 7973.6 4677.9 4213.3 27234.0 29654.6 10958.9 25832.4 20695.4 27527.3 82677.8 137588.1 84576.7 195401.4 50300.9 49717.2 51772.3 549.1 450.5 358.3 918.2
JapaneseVowels 68.0 23.2 511.6 4572.5 130.6 220.0 464.5 1468.1 654.1 375.3 422.4 105.1 217.7 313.5 370.0 288.0 786.6 634.8 2156.2 2337.2 988.1 2354.9 2845.0 2385.2 11809.2 23292.2 7908.2 12345.3 8072.1 8403.0 8069.3 83.1 263.0 178.9 190.3
ibm-employee-attrition 11.4 4.4 35.2 841.8 6.6 17.8 27.2 44.6 30.5 20.2 28.4 16.6 27.3 30.9 53.0 26.2 60.8 59.3 634.9 615.6 311.2 541.1 602.7 529.5 1544.5 2578.5 1444.5 1962.3 1199.1 1067.7 1165.2 11.4 276.5 247.6 229.7
first-order-theorem-proving 260.6 123.0 1076.6 14412.8 189.2 564.0 376.0 377.1 451.6 581.2 696.8 494.5 1209.6 1387.6 16399.4 1393.4 507.1 753.1 14029.4 14742.7 2543.1 14821.9 22683.2 15135.4 54801.6 149905.1 60861.9 113144.1 63868.9 70784.0 67082.8 137.9 970.4 1264.2 480.2
user-knowledge 1.2 1.0 9.4 445.3 3.2 2.4 7.0 6.0 3.6 9.6 8.0 1.6 5.2 8.0 20.8 1.2 75.1 22.2 249.5 182.0 136.8 174.6 186.5 168.1 250.2 284.3 205.6 496.7 139.2 153.2 137.8 1.4 279.2 285.1 171.8
online-shoppers-intention 59.0 35.7 932.5 3352.5 70.6 357.1 345.9 1500.2 1245.3 392.7 432.3 89.4 152.9 191.9 1065.0 193.9 284.6 321.6 2719.6 3234.2 1051.6 2275.9 2352.1 2280.6 8023.8 15974.0 7939.9 11193.9 5435.3 5324.7 5426.7 57.8 287.1 238.8 337.2
kc1 11.0 4.8 43.4 1018.0 4.2 17.2 28.8 145.8 145.9 21.4 29.4 12.6 27.2 28.6 62.2 29.0 71.6 81.2 548.6 559.3 283.7 444.9 551.6 467.4 2477.4 4628.3 1163.0 1747.6 1503.6 1808.2 1488.7 9.4 255.5 227.4 173.4
thoracic-surgery 1.4 1.0 7.0 462.5 2.0 2.6 2.8 9.0 17.9 3.8 7.6 2.2 3.2 4.0 9.3 4.0 20.8 16.2 222.5 200.9 134.8 199.3 253.7 211.5 348.6 415.7 265.5 358.0 192.2 200.4 200.9 1.8 401.1 686.9 164.3
UCI_churn 22.7 6.2 77.9 1875.7 11.6 46.2 39.6 99.2 70.3 47.2 64.7 25.6 41.8 45.8 106.6 46.0 110.4 118.5 688.5 657.5 342.2 705.0 981.7 687.2 2986.1 4644.7 2223.0 3048.3 2640.1 2433.5 2638.2 24.7 263.7 211.4 134.5
arsenic-female-bladder 1.0 1.0 7.0 529.0 2.0 2.4 2.4 8.6 9.1 9.4 9.5 2.4 4.8 5.0 11.3 4.6 32.1 19.2 249.8 199.8 154.9 196.6 209.8 186.1 287.5 321.7 217.3 314.1 153.1 152.8 157.3 1.4 260.3 95.6 113.2
okcupid_stem 677.7 170.8 33141.2 29727.6 1770.0 13170.7 9867.2 9610.6 9452.7 12852.7 20467.5 3914.3 5342.9 8847.6 319856.7 11635.3 1804.5 2004.5 60835.2 60473.4 5368.4 43825.6 39398.3 46729.0 223807.9 215394.0 349760.3 480048.2 98005.7 102574.4 99067.5 278.4 634.3 702.2 673.9
ecoli 0.8 1.2 7.8 396.5 3.6 2.2 6.6 17.5 9.4 0.8 7.2 1.6 4.8 7.6 17.6 8.6 75.2 21.7 243.7 179.0 137.0 172.7 177.2 179.7 252.0 292.5 206.6 506.3 133.2 139.4 129.0 1.4 288.7 291.4 124.4
pc4 8.6 4.0 29.2 887.2 6.8 15.4 22.0 100.6 298.3 64.6 170.8 20.4 153.4 106.4 198.1 201.8 81.9 88.1 936.3 1917.4 1346.5 1291.6 1792.9 906.4 5024.9 14518.1 2946.5 2456.0 984.5 1078.9 975.6 10.1 288.8 1073.6 216.6
bank-marketing 22.6 5.9 140.6 1419.9 17.8 71.1 70.0 285.6 250.8 80.3 104.3 36.4 70.0 84.8 178.2 70.3 149.1 164.4 1097.8 1163.9 311.1 875.8 746.8 857.2 3243.3 6964.4 3621.8 4904.8 2288.0 2016.9 2198.1 22.4 269.0 216.1 254.8
Diabetes-130-Hospitals_(Fairlearn) 561.7 142.8 106332.6 26681.7 2729.1 23801.1 20950.1 120267.6 50427.3 26071.4 24390.4 1099.3 2066.7 4550.6 3751167.6 4560.9 5776.2 7213.0 22111.3 22994.3 6756.0 20215.9 14156.5 20308.6 102571.3 177673.0 115777.3 160993.5 51156.7 53470.3 50926.5 577.7 549.8 370.3 1097.1
Otto-Group-Product-Classification-Challenge 911.5 210.1 12390.7 27696.4 1530.4 12848.0 38053.7 53203.7 51895.8 50469.0 57542.1 2118.4 21850.3 58771.5 1403163.3 58822.2 10757.4 11078.1 66475.9 73756.5 12228.7 23946.1 26298.5 23410.6 246684.7 1259334.9 199213.6 1077151.2 94282.3 89978.2 94552.6 911.8 1876.1 3178.6 1664.0
eucalyptus 16.0 4.7 88.7 1424.1 9.8 52.1 45.9 137.7 72.9 63.0 92.6 26.4 47.2 71.4 692.7 70.0 131.1 137.0 866.6 735.0 282.9 590.3 563.3 560.5 2452.2 3839.1 2338.2 3219.2 1578.1 1598.7 1574.5 17.2 411.8 844.7 247.9
pendigits 24.0 8.0 463.6 1996.3 36.1 228.1 217.9 816.1 579.0 194.6 311.8 51.2 146.3 161.4 205.6 173.0 284.7 302.0 883.3 821.5 505.2 865.2 905.5 879.9 5336.6 17508.9 4429.2 9634.1 2569.5 2244.1 2550.7 22.2 250.7 181.8 191.0
pc3 18.2 3.4 26.0 1001.1 5.4 20.5 16.8 35.7 28.5 23.1 34.3 24.2 52.8 57.2 59.8 49.2 97.3 97.0 500.7 456.7 258.8 444.8 475.9 444.1 2385.4 5365.0 1605.7 2869.2 1425.5 1567.1 1397.1 13.8 333.1 355.1 287.8
page-blocks-bin 24.2 4.8 125.4 2337.0 19.8 60.8 114.2 263.1 166.3 104.8 112.7 21.4 42.5 55.6 107.4 53.4 186.8 132.0 521.9 517.8 274.6 477.5 617.5 492.6 2466.9 4182.5 1742.8 2525.9 2289.6 1803.9 2281.3 19.0 255.7 170.1 148.0
optdigits 30.3 7.8 239.6 1410.9 18.8 111.6 136.8 333.9 372.0 111.4 159.5 60.6 184.4 266.6 239.9 247.2 303.9 327.1 879.5 874.6 390.8 995.2 1176.1 969.3 6458.1 21344.9 5498.7 12072.3 3027.8 2535.2 2986.0 28.0 244.0 240.0 281.7
mfeat-karhunen 64.0 7.8 51.2 2307.1 11.0 74.7 81.8 121.1 104.4 86.5 105.6 61.2 118.0 124.3 160.7 161.4 202.4 217.2 913.3 847.1 368.1 809.2 2082.6 827.9 6111.3 13804.2 4229.5 6882.7 7206.2 5055.5 7092.1 47.6 261.8 177.3 240.0
[…]